# Learnable Infinite Taylor Gaussian for Dynamic View Rendering

### v1

@@ 1 Introduction
Figure 1: Comparison of Methods based on the mean values of SSIM and PSNR on the Neural 3D Video Dataset [19].
Recently, 3D Gaussian Splatting (3DGS) has achieved groundbreaking progress in dynamic scene reconstruction, especially with the adoption of tile-based rasterization techniques as a replacement for traditional volumetric rendering methods [15, 18, 23, 26, 24]. This innovation has garnered substantial attention within the academic community. Many researchers have started to leverage 3DGS for 4D scene reconstruction [38, 39, 13, 27], aiming to accurately capture and model the evolving 3D structure and appearance of scenes over time, enabling novel view synthesis at arbitrary time points. Although modeling static scenes has seen significant advances [15, 5, 21, 43, 7, 6], dynamic scene reconstruction remains challenging due to factors such as the complexity of object motion, topological changes, and spatial or temporal sparsity in observations [27, 9, 34, 23, 18]. These issues make accurate reconstruction of dynamic scenes a technical challenge, requiring ongoing research and innovation to overcome.
@@
To address the challenge of capturing the temporal evolution of Gaussian motion attributes, we propose an innovative 3DGS framework. As shown in Figure  2, this framework deeply explores the mathematical principles behind Gaussian point motion and analyzes their trajectories to accurately simulate complex motion dynamics. Specifically, we introduce a learnable infinite Taylor series to model the motion trajectories of Gaussian points in dynamic scenes. By tracking the evolution of Gaussian points over time, we can precisely capture key attributes (such as position, opacity, and scale) at each time step. This approach not only provides a solid mathematical foundation for 3D reconstruction and view synthesis but also offers a novel perspective for dynamic scene modeling.
@@ 4.1 Fundamental Theory of Taylor Formula
$\begin{split}f(x)=&f(x_{0})+f^{{}^{\prime}}(x_{0})(x-x_{0})+\frac{1}{2!}f^{{}^{\prime\prime}}(x_{0})(x-x_{0})^{2}\\ &+\frac{1}{3!}f^{(3)}(x_{0})(x-x_{0})^{2}+\dots\\ &+\frac{1}{n!}f^{(n)}(x_{0})(x-x_{0})^{n}+R_{n}(x)\end{split}$ (8)
@@ 4.3 Taylor Expansion of Transformation Field Modeling
$\bm{p}_{i}(t)=\sum_{k=0}^{n}\frac{1}{k!}f_{p}^{(k)}\left(t_{\tau}\right)(t-t_{\tau})^{k},$ (12)
@@
$\bm{s}_{i}(t)=\sum_{k=0}^{m}\frac{1}{k!}f_{s}^{(k)}\left(t_{\tau}\right)(t-t_{\tau})^{k},$ (13)
@@
$\bm{q}_{i}(t)=\sum_{k=0}^{l}\frac{1}{k!}f_{q}^{(k)}\left(t_{\tau}\right)(t-t_{\tau})^{k},$ (14)
@@ 4.4 Peano Remainder of Transformation Fields Modeling
$w_{ij}=\frac{\hat{w}_{ij}}{\sum_{j\in\mathcal{N}}\hat{w}_{ij}}\text{, where }\hat{w}_{ij}=\exp\left(-\frac{d_{ij}^{2}}{2r_{j}^{2}}\right),$ (16)
@@
Figure 3: Comparison of novel view rendering on the N3DV dataset, with problem regions highlighted in boxes. More results are provided in the supplementary.
@@ 5 Experiments
Table 1: Comparison of methods in novel view rendering based on the N3DV dataset.
Method Cook Spinach Sear Steak Flame Steak
[…]
MixVoxels [37] 31.39 0.931 0.113 30.85 0.940 0.103 30.15 0.938 0.108
K-Planes [8] 31.23 0.926 0.114 30.28 0.937 0.104 31.49 0.940 0.102
HexPlane [2] 31.05 0.928 0.114 30.00 0.939 0.105 30.42 0.939 0.104
HyperReel [1] 31.77 0.932 0.090 31.88 0.942 0.080 31.48 0.939 0.083
NeRFPlayer [35] 30.58 0.929 0.113 29.13 0.908 0.138 31.93 0.950 0.088
StreamRF [18] 30.89 0.914 0.162 31.60 0.925 0.147 31.37 0.923 0.152
SWinGS [34] 31.96 0.946 0.094 32.21 0.950 0.092 32.18 0.953 0.087
D3DGS [39] 20.53 0.881 0.153 25.02 0.944 0.072 23.02 0.919 0.113
4DGS [38] 28.12 0.940 0.038 29.07 0.957 0.028 25.04 0.918 0.079
FSGS [44] 29.60 0.919 0.115 29.87 0.945 0.116 29.42 0.943 0.113
Ours 30.17 0.945 0.073 38.43 0.985 0.030 37.64 0.984 0.029
Method Flame Salmon Coffee Martini Cut Roast Beef
[…]
MixVoxels [37] 29.50 0.898 0.163 29.25 0.901 0.147 31.38 0.928 0.111
K-Planes [8] 29.58 0.901 0.132 29.30 0.900 0.134 31.87 0.928 0.114
HexPlane [2] 29.23 0.905 0.088 28.45 0.891 0.149 30.83 0.927 0.115
HyperReel [1] 28.26 0.941 0.136 28.65 0.897 0.129 32.25 0.936 0.086
NeRFPlayer [35] 31.65 0.940 0.098 31.53 0.951 0.085 29.35 0.908 0.144
StreamRF [18] 28.69 0.872 0.229 28.13 0.873 0.219 30.75 0.917 0.154
SWinGS [34] 29.25 0.925 0.100 29.16 0.921 0.105 31.84 0.945 0.099
D3DGS [39] 23.03 0.863 0.153 23.42 0.865 0.138 22.35 0.907 0.125
4DGS [38] 27.90 0.916 0.063 28.18 0.924 0.049 29.71 0.944 0.033
FSGS [44] 28.06 0.926 0.102 26.57 0.907 0.169 28.46 0.913 0.122
Ours 31.85 0.972 0.075 32.52 0.974 0.074 36.96 0.984 0.035
@@ 5.1 Implementation Details
During the experimental phase, we optimize processing efficiency by downsampling the image resolution by a factor of 2. By using the Adam optimizer with an adaptive learning rate, we are able to accelerate training and improve model performance. In our setup, a single NVIDIA RTX 4090 GPU provides sufficient computational power to handle these downsampled images.
@@ 5.2 Datasets and Metrics
This paper uses two real-world datasets: Neural 3D Video (N3DV) [19] and MeetRoom [18]. The N3DV dataset is captured using a 21-camera multi-view system, while the MeetRoom dataset is recorded with a multi-view setup consisting of 13 Azure Kinect cameras. Specifically, the N3DV dataset includes six distinct scenes: Coffee Martini, Cook Spinach, Cut Roasted Beef, Flame Salmon, Flame Steak, and Sear Steak. Each scene contains 300 frames, featuring extended durations and diverse actions, with some scenes incorporating multiple moving objects. In the collection of the MeetRoom dataset benchmark, the system synchronizes all camera shutters based on external devices. Compared to other commonly used datasets, such as LLFF [28] and N3DV [19], the MeetRoom dataset has more sparsely aligned views.
@@ 5.3 Novel View Rendering on N3DV
As shown in Table 1, our method outperforms others overall, achieving relatively strong performance, with the best PSNR and SSIM scores on each sub-dataset, except for Cook Spinach. The Mixvoxel algorithm [37] converts point cloud data into a regular voxel grid, simplifying the data processing pipeline. This voxel representation allows for efficient feature extraction. However, during the voxelization process, some fine-grained details may be lost due to the conversion of point cloud data into a regular grid. As a result, the algorithm’s performance may be affected in complex scenarios. Moreover, the performance of the Mixvoxel algorithm is sensitive to voxel parameters, such as voxel size and resolution. Different parameter settings can lead to significant fluctuations in performance, requiring careful adjustment and optimization. This adds complexity and challenges in applying the algorithm to practical tasks.
@@
Table 2: Methods comparison in MeetRoom dataset. More resutls will be illustrated in the supplementary.
Method Discussion Vrheadset
[…]
STG [22] 25.53 0.881 0.193 23.13 0.847 0.244
D3DGS [39] 32.52 0.956 0.041 32.10 0.962 0.037
4DGS [38] 30.61 0.944 0.160 32.12 0.946 0.044
FSGS [44] 26.82 0.921 0.198 26.27 0.903 0.257
Ours 36.77 0.977 0.014 34.81 0.980 0.042
Method Trimming Stepin
[…]
STG [22] 22.54 0.819 0.321 24.14 0.857 0.234
D3DGS [39] 30.55 0.946 0.061 36.19 0.968 0.034
4DGS [38] 31.18 0.934 0.065 36.16 0.961 0.036
FSGS [44] 25.68 0.901 0.232 31.48 0.947 0.196
Ours 34.28 0.975 0.050 36.35 0.977 0.050
@@
5.4 Novel View Rendering on MeetRoom
As shown in Table 2, the proposed method yields significantly more accurate and robust results. For example, in the Discussion sequence, the PSNR of the proposed method is 36.77, whereas the state-of-the-art methods STG [22], D3DGS [39], 4DGS [38], and FSGS [44] achieve PSNR values of 25.53, 32.52, 30.61, and 26.82, respectively. This results in improvements of approximately $\mathbf{44.02\%}$, $\mathbf{13.07\%}$, $\mathbf{20.12\%}$, and $\mathbf{37.10\%}$ over these methods. Similar trends are observed in other sequences, including Vrheadset, Trimming, and Stepin. Compared to STG [22] and 4DGS [38], the D3DGS [39] method demonstrates more robust performance across sequences, particularly in the Stepin sequence, where it achieves the best LPIPS score and also delivers competitive PSNR and SSIM results.
@@ 5.5 Ablation Study
Table 3: Ablation studies of module performance in novel view rendering using the Cut Roast Beef dataset.
@@
w/o Time-opacity 32.10 0.969 0.046
w/o Time-motion 29.80 0.961 0.051
w/o Time-rotation 32.94 0.972 0.053
w/o Time-scale 27.13 0.935 0.052
w/o Peano remainder 28.15 0.944 0.049
Ours Full 36.96 0.984 0.035
@@ Appendix A Novel View Rendering
In the evaluation section, we have designed a series of comprehensive experiments to assess the performance of our method. Here, we further provide a set of quantitative and qualitative results to more thoroughly validate the effectiveness of our approach.
@@ A.1 Qualitative Analysis of Details
In the quantitative analysis of novel view rendering algorithms, we focused on several key evaluation metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM). We Learned Perceptual Image Patch Similarity (LPIPS) [42]. These metrics help us quantify the details and overall quality of the reconstructed images. By comparing these metrics, we can more accurately assess the performance differences between different algorithms. To gain a more comprehensive understanding, we also incorporated qualitative analysis, examining the detailed performance of various algorithms in the reconstructed images, leading to a deeper evaluation. Through visual presentation, we can further assess the strengths and weaknesses of the algorithms, ensuring a multidimensional and comprehensive understanding of their performance.
@@
As shown in Figure 4, we can see that our algorithm demonstrates superior performance in detail reconstruction compared to others. However, D3DGS and 4DGS face challenges such as blurriness and distortion during the reconstruction process. We provide a detailed explanation of each row in Figure 4:
First Row: D3DGS and 4DGS struggle with blurry reconstructions in areas with fine details, accompanied by significant noise. In scenes with glass occlusion, the recovery of intricate details is insufficient, further diminishing reconstruction quality.
Second Row: Texture reconstruction by D3DGS and 4DGS is subpar, evident in blurry black dots and the incomplete restoration of the green bottle. In complex scenarios with multiple objects, such as the region highlighted in the green box, the reconstruction between bottles reveals significant flaws.
Third Row: Both methods perform poorly in occluded scenes. For instance, glass occlusion hampers the accurate reconstruction of scenes behind the glass. Similarly, bottle occlusion and glass refraction cause deformation and blurriness in the reconstruction of the orange object.
Fourth Row: D3DGS and 4DGS exhibit color distortion, particularly under indoor lighting conditions. Additionally, certain areas, such as the edges of the pan and text, appear blurry. The images lack contrast; for example, in the region marked by the red box, reflective textures and hair details are inaccurately reconstructed.
@@ A.2 Qualitative Analysis of Ablation Experiments
In our ablation experiments on the Cut Roasted Beef class in the N3DV dataset, we conducted an in-depth qualitative analysis to evaluate the impact of ablating different modules on the performance of reconstructed images and the representation of fine details. By systematically comparing images reconstructed after the ablation of various modules, we were able to uncover their respective strengths and limitations in handling complex scenes.
First, significant differences were observed in rendering quality across the ablations of different modules. Ablating specific modules reduced the ability to capture geometric details of objects, as shown in Figure 5, such as surface textures and edge contours. For example, as highlighted by the blue bounding box, both w/o Peano remainder and w/o Time-scale failed to accurately depict geometric details, leading to missing geometric information. Similarly, as shown in the green bounding box, w/o Peano remainder and w/o Time-motion exhibited poor performance in reconstructing surface textures, producing artifacts such as shadowing and black linear streaks. Additionally, w/o Time-scale demonstrated a weaker capability to capture edge contours, resulting in blurred or muddled details during reconstruction. In other cases, module ablations caused noticeable noise or distortions in specific details. For instance, as illustrated in the red bounding box, w/o Time-opacity and w/o Time-rotation introduced significant noise when reconstructing fine details such as hair compared to the original images. These differences were particularly pronounced when processing Cut Roasted Beef samples with rich geometric features, highlighting the critical role of these modules in maintaining reconstruction fidelity.
Figure 5: Cut Roasted Beef Novel View Rendering on the N3DV Dataset: Qualitative Analysis of Ablation Experiments - Comparison of Reconstruction Quality and Detail Representation with Module Ablations.
Table 4: Methods comparison in Technicolor dataset.
Method Birthday Painter
[…]
D3DGS[39] 33.81 0.965 0.014 37.38 0.957 0.036
STG[22] 33.87 0.951 0.038 37.30 0.928 0.095
FSGS[44] 26.26 0.920 0.068 34.36 0.958 0.063
4DGS[38] 21.94 0.902 0.071 28.61 0.940 0.058
Ours 34.72 0.988 0.013 38.37 0.985 0.022
Furthermore, we examined the ability of different module ablations to handle lighting variations. As depicted in the yellow bounding box, w/o Time-motion, w/o Time-opacity, and w/o Time-rotation often reconstructed converging light rays as divergent, leading to distorted lighting effects that compromised overall visual quality. This underscores the importance of these modules in accurately capturing lighting dynamics in the scene.
@@
A.3 Comprehensive Evaluation from a Quantitative Analysis Perspective
Table 5: Comprehensive Evaluation Metrics of Different Algorithms on the N3DV Dataset.
Method Mean
[…]
MixVoxels [37] 30.42 0.923 0.124
K-Planes [8] 30.63 0.922 0.117
HexPlane [2] 30.00 0.922 0.113
HyperReel [1] 30.72 0.931 0.101
NeRFPlayer [35] 30.70 0.931 0.111
StreamRF [18] 30.24 0.904 0.177
SWinGS [34] 31.10 0.940 0.096
D3DGS [39] 22.90 0.897 0.126
4DGS [38] 28.00 0.933 0.048
FSGS [44] 28.66 0.926 0.123
Ours 34.60 0.974 0.053
We conduct a comprehensive evaluation of the performance of different algorithms on the N3DV dataset, performing an comparative analysis based on the mean values of each evaluation metric. As listed in Table 5, we assess the performance of each algorithm on the N3DV dataset and computed the mean values of PSNR, SSIM, and LPIPS across all scenes. The proposed approach enables us to accurately reflect the overall performance of each algorithm across all scenarios, ensuring the stability and comparability of the evaluation results.
Furthermore, our algorithm demonstrates significant advantages across these metrics, particularly in the PSNR and SSIM composites, as illustrated in Figure 1. Our method consistently outperforms other benchmark algorithms, showcasing its superior ability to preserve detail and structural fidelity across all scenes. Compared to state-of-the-art methods, our algorithm is more effective at suppressing noise and improving image quality, especially in complex dynamic scenes.
Table 6: Comprehensive Evaluation Metrics of Different Algorithms on the MeetRoom Dataset.
Method Mean
[…]
STG [22] 23.83 0.851 0.248
D3DGS [39] 32.84 0.958 0.043
4DGS [38] 32.52 0.946 0.076
FSGS [44] 27.56 0.918 0.221
Ours 35.55 0.977 0.039
To further validate the superiority of our algorithm, we conducted the same evaluation on the MeetRoom dataset and computed the mean values for each evaluation metric, as shown in Table 6. The results from this dataset further reinforce the efficiency and robustness of our algorithm across diverse scenarios. Specifically, our method continues to excel across all metrics, particularly in terms of PSNR and SSIM, further confirming its remarkable ability to enhance detail recovery and image quality. Similar to the results on the N3DV dataset, the outstanding performance of our algorithm on the MeetRoom dataset reaffirms its superiority in handling complex scenes.
In summary, the proposed method not only achieves exceptional performance in traditional quality assessment metrics but also demonstrates remarkable robustness and adaptability in comprehensive evaluations, consistently delivering superior performance across different datasets.

### v2

@@ 1 Introduction
Recently, 3D Gaussian Splatting (3DGS) has achieved groundbreaking progress in dynamic scene reconstruction [44, 39, 41], especially with the adoption of tile-based rasterization techniques as a replacement for traditional volumetric rendering methods [15, 18, 23, 26, 24, 11]. This innovation has garnered substantial attention within the academic community. Many researchers have started to leverage 3DGS for 4D scene reconstruction [39, 40, 13, 27], aiming to accurately capture and model the evolving 3D structure and appearance of scenes over time, enabling novel view synthesis at arbitrary time points. Although modeling static scenes has seen significant advances [15, 5, 21, 45, 6], dynamic scene reconstruction remains challenging due to factors such as the complexity of object motion, topological changes, and spatial or temporal sparsity in observations [27, 8, 35, 23, 18]. These issues make accurate reconstruction of dynamic scenes a technical challenge, requiring ongoing research and innovation to overcome.
@@
To address the challenge of capturing the temporal evolution of Gaussian motion attributes, we propose an innovative 3DGS framework. As shown in Figure  1, this framework deeply explores the mathematical principles behind Gaussian point motion and analyzes their trajectories to accurately simulate complex motion dynamics. Specifically, we introduce a learnable infinite Taylor series to model the motion trajectories of Gaussian points in dynamic scenes. By tracking the evolution of Gaussian points over time, we can precisely capture key attributes (such as position, opacity, and scale) at each time step. This approach not only provides a solid mathematical foundation for 3D reconstruction and view synthesis but also offers a novel perspective for dynamic scene modeling.
@@ 4.1 Fundamental Theory of Taylor Formula
$\begin{split}f(x)=&f(x_{0})+f^{{}^{\prime}}(x_{0})(x-x_{0})+\frac{1}{2!}f^{{}^{\prime\prime}}(x_{0})(x-x_{0})^{2}\\ &+\frac{1}{3!}f^{(3)}(x_{0})(x-x_{0})^{3}+\dots\\ &+\frac{1}{n!}f^{(n)}(x_{0})(x-x_{0})^{n}+R_{n}(x)\end{split}$ (8)
@@ 4.3 Taylor Expansion of Transformation Field Modeling
$\bm{p}_{i}(t)=\sum_{k=0}^{n}\frac{1}{k!}f_{p}^{(k)}\left(t_{\tau}\right)(t-t_{\tau})^{k}$ (12)
@@
$\bm{s}_{i}(t)=\sum_{k=0}^{m}\frac{1}{k!}f_{s}^{(k)}\left(t_{\tau}\right)(t-t_{\tau})^{k}$ (13)
@@
$\bm{q}_{i}(t)=\sum_{k=0}^{l}\frac{1}{k!}f_{q}^{(k)}\left(t_{\tau}\right)(t-t_{\tau})^{k}$ (14)
@@ 4.4 Peano Remainder of Transformation Fields Modeling
$w_{ij}=\frac{\hat{w}_{ij}}{\sum_{j\in\mathcal{N}}\hat{w}_{ij}}\text{, where }\hat{w}_{ij}=\exp\left(-\frac{d_{ij}^{2}}{2r_{j}^{2}}\right)$ (16)
@@
Figure 2: Comparison of novel view rendering on the N3DV dataset, with problem regions highlighted in boxes. More results can be found in the supplementary material and on our project website.
@@ 5.1 Implementation Details
During the experimental phase, we use COLMAP [33] to reconstruct the geometric structure of 3D scenes, including point clouds and camera poses, providing each model with high-quality initial point cloud data. By leveraging the Adam optimizer with an adaptive learning rate and a single NVIDIA RTX 4090 GPU, we effectively accelerate training and enhance model performance.
@@ 5.2 Datasets and Metrics
This study utilizes two real-world datasets: Neural 3D Video (N3DV) [19] and the Technicolor Light Field Dataset [32]. The N3DV dataset is captured using a multi-view system consisting of 21 cameras, while the Technicolor dataset records video sequences using a $4\times 4$ array of 16 cameras. These cameras are precisely synchronized in time and can capture high-resolution images with a spatial resolution of up to $2048\times 1088$ pixels. Specifically, we select four different scenes from the N3DV dataset: Cook Spinach, Cut Roasted Beef, Flame Steak, and Sear Steak. Each scene consists of 300 frames, featuring extended durations and diverse motions, with some scenes containing multiple moving objects. For the Technicolor Light Field Dataset, we choose four distinct scenes: Birthday, Painter, Train, and Fatma. These scenes not only enhance the dataset’s diversity but also provide a comprehensive testing environment for model evaluation.
Table 1: Comparison of methods in novel view rendering based on the N3DV dataset. Best results are highlighted in bold.
Method Cook Spinach Sear Steak Flame Steak Cut Roast Beef
[…]
MixVoxels [38] 31.39 0.931 0.113 30.85 0.940 0.103 30.15 0.938 0.108 31.38 0.928 0.111
K-Planes [7] 31.23 0.926 0.114 30.28 0.937 0.104 31.49 0.940 0.102 31.87 0.928 0.114
HexPlane [2] 31.05 0.928 0.114 30.00 0.939 0.105 30.42 0.939 0.104 30.83 0.927 0.115
HyperReel [1] 31.77 0.932 0.090 31.88 0.942 0.080 31.48 0.939 0.083 32.25 0.936 0.086
NeRFPlayer [36] 30.58 0.929 0.113 29.13 0.908 0.138 31.93 0.950 0.088 29.35 0.908 0.144
StreamRF [18] 30.89 0.914 0.162 31.60 0.925 0.147 31.37 0.923 0.152 30.75 0.917 0.154
SWinGS [35] 31.96 0.946 0.094 32.21 0.950 0.092 32.18 0.953 0.087 31.84 0.945 0.099
D3DGS [40] 20.53 0.881 0.153 25.02 0.944 0.072 23.02 0.919 0.113 22.35 0.907 0.125
4DGS [39] 28.12 0.940 0.038 29.07 0.957 0.028 25.04 0.918 0.079 29.71 0.944 0.033
SCGS [12] 17.20 0.734 0.232 28.77 0.951 0.056 23.49 0.902 0.104 6.29 0.007 0.683
Ours 32.59 0.966 0.054 33.12 0.973 0.049 33.34 0.971 0.052 33.06 0.969 0.055
Table 2: Methods comparison on the Technicolor dataset. Best results are highlighted in bold.
Method Birthday Painter Train Fatma
[…]
D3DGS[40] 33.81 0.965 0.014 37.38 0.957 0.036 - - - 38.40 0.911 0.093
STG[22] 33.87 0.951 0.038 37.30 0.928 0.095 33.36 0.948 0.036 37.28 0.906 0.155
FSGS[46] 26.26 0.920 0.068 34.36 0.958 0.063 30.39 0.965 0.032 27.62 0.825 0.276
4DGS[39] 21.94 0.902 0.071 28.61 0.940 0.058 22.36 0.878 0.124 23.42 0.763 0.236
Ours 34.72 0.988 0.013 38.37 0.985 0.022 35.30 0.990 0.008 38.91 0.945 0.071
@@ 5.3 Novel View Rendering on N3DV
As shown in Table 1, our method outperforms others overall, achieving relatively strong performance, with the best PSNR and SSIM scores on each sub-dataset. The Mixvoxel algorithm [38] converts point cloud data into a regular voxel grid, simplifying the data processing pipeline. This voxel representation allows for efficient feature extraction. However, during the voxelization process, some fine-grained details may be lost due to the conversion of point cloud data into a regular grid. As a result, the algorithm’s performance may be affected in complex scenarios. Moreover, the performance of the Mixvoxel algorithm is sensitive to voxel parameters, such as voxel size and resolution. Different parameter settings can lead to significant fluctuations in performance, requiring careful adjustment and optimization. This adds complexity and challenges in applying the algorithm to practical tasks.
@@
5.4 Novel View Rendering on Technicolor Dataset
Here, we conduct a detailed comparison against state-of-the-art approaches to further validate the superiority of our method. By analyzing multiple sequences within the dataset, we highlight the consistency and robustness of our approach in handling diverse and complex scenes. The following quantitative results illustrate the substantial improvements achieved by our method over existing techniques.
As shown in Table 2, the proposed method yields significantly more accurate and robust results. For example, in the Birthday sequence, the PSNR of the proposed method is 34.72, whereas the state-of-the-art methods 4DGS [39], FSGS [46], D3DGS [40], and STG [22] achieve PSNR values of 21.94, 26.26, 33.81, and 33.87, respectively. This results in improvements of approximately $\mathbf{58.25\%}$, $\mathbf{32.22\%}$, $\mathbf{2.69\%}$, and $\mathbf{2.51\%}$ over these methods. Similar trends are observed in other sequences, including Painter, Train, and Fatma. Compared to STG [22] and 4DGS [39], the D3DGS [40] method demonstrates more robust performance across sequences, particularly in the Painter, Fatma sequence, where it achieves the best PSNR score and also delivers competitive LPIPS and SSIM results.
@@ 5.5 Ablation Study
Table 3: Ablation study on the N3DV dataset. best results are highlighted in bold.
@@
w/o Time-opacity 31.17 0.952 0.096
w/o Time-motion 29.24 0.920 0.154
w/o Time-rotation 31.21 0.953 0.103
w/o Time-scale 31.40 0.953 0.097
w/o Peano remainder 31.51 0.935 0.103
Ours Full 33.03 0.970 0.052
@@ 6 Discussion and Conclusion
7 Acknowledgments
The research/project is supported by the National Research Foundation (NRF) Singapore, under its NRF-Investigatorship Programme (Award ID. NRF-NRFI09-0008), the National Natural Science Foundation of China (No.62072388), Fujian Provincial Science and Technology Major Project (No.2024HZ022003), Jiangxi Provincial Natural Science Foundation Key Project (No.20244BAB28039), Xiamen Public Technology Service Platform (No.3502Z20231043), and Fujian Sunshine Charity Public Welfare Foundation.
@@ Appendix A Novel View Rendering
In the evaluation section, we have designed a series of comprehensive experiments to assess the performance of our method. Here, we present a set of visual results to further validate the effectiveness of our approach more thoroughly.
@@ A.1 Qualitative Analysis of Details
In the quantitative analysis of novel view rendering algorithms, we focused on several key evaluation metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) and Perceptual Image Patch Similarity (LPIPS) [43]. These metrics help us quantify the details and overall quality of the reconstructed images. By comparing these metrics, we can more accurately assess the performance differences between different algorithms. To gain a more comprehensive understanding, we also incorporated qualitative analysis, examining the detailed performance of various algorithms in the reconstructed images, leading to a deeper evaluation. Through visual presentation, we can further assess the strengths and weaknesses of the algorithms, ensuring a multidimensional and comprehensive understanding of their performance.
@@
As shown in Figure 3, we can see that our algorithm demonstrates superior performance in detail reconstruction compared to others. However, D3DGS and 4DGS face challenges such as artifacts and distortions during the reconstruction process. We provide a detailed explanation of each row in Figure 3:
First Row: Overall, the texture reconstruction quality of D3DGS and 4DGS is below the standard. Additionally, as seen in the red box (curtain reconstruction) and the blue box (wall reconstruction), the detail reconstruction performance of D3DGS and 4DGS is also poor.
Second Row: The overall reconstruction performance of D3DGS is poor. Both 4DGS and D3DGS exhibit issues in detail reconstruction, such as blurred shadows in the yellow box, significant reflections and artifacts on the leather stool in the red box, and additional artifacts appearing in the blue box for D3DGS.
Third Row: D3DGS performs poorly in both overall reconstruction quality and detail representation. 4DGS also has some issues in detail reconstruction, such as unexplained black spots above the white bottle in the green box and unexplained light appearing on the left side of the blue box.
Fourth Row: Both D3DGS and 4DGS exhibit color distortions. Additionally, shadows appear in certain areas (e.g., red, yellow, and blue boxes), and the image in the blue box lacks contrast. There are also extraneous elements in the green box of D3DGS.
@@ A.2 Qualitative Analysis of Ablation Experiments
In our ablation experiments on the Sear Steak class in the N3DV dataset, we conducted an in-depth qualitative analysis to evaluate the impact of ablating different modules on the performance of reconstructed images and the representation of fine details. By systematically comparing images reconstructed after the ablation of various modules, we were able to uncover their respective strengths and limitations in handling complex scenes.
First, significant differences were observed in rendering quality across the ablations of different modules. Ablating specific modules reduced the ability to capture geometric details of objects, as shown in Figure 4, such as surface textures and edge contours. For example, as highlighted by the blue bounding box, both w/o Peano remainder and w/o Time-opacity failed to accurately capture geometric details, leading to missing geometric information. Similarly, as shown in the green bounding box, w/o Peano remainder, w/o Time-opacity, and w/o Time-scale exhibited poor performance in reconstructing surface textures, producing artifacts such as shadowing and linear streaks. Additionally, w/o Peano remainder and w/o Time-opacity demonstrated a weaker capability in capturing edge contours, resulting in blurred or muddled details during reconstruction.
In other cases, module ablations introduced noticeable noise or over-smoothing in specific details. For instance, as illustrated in the red bounding box, w/o Time-motion and w/o Time-rotation introduced significant noise when reconstructing fine details compared to the original images. These differences were particularly pronounced when processing Sear Steak samples with rich geometric features, highlighting the critical role of these modules in maintaining reconstruction fidelity. Furthermore, we evaluated the impact of different module ablations on handling complex scenes. As shown in the yellow bounding box, the reconstruction quality of w/o Time-motion, w/o Time-rotation, and w/o Time-scale was relatively blurry, with increased noise and excessive smoothing, ultimately degrading the overall visual quality. This underscores the importance of these modules in accurately capturing fine details in complex scenes.
Figure 4: Sear Steak Novel View Rendering on the N3DV Dataset: Qualitative Analysis of Ablation Experiments - Comparison of Reconstruction Quality and Detail Representation with Module Ablations.